The equation below describes an economy's macroeconomic structure measured in millions
C =0.5(1 - t)Y
t =0.20
I =800 - 30i
G =700
L=0.3Y - 60i
M/P =600
a) for this economy, use the relevant set of equations to derive and graph the IS-LM relations
1
Expert's answer
2021-08-30T17:49:59-0400

"Solution"

IS CURVE

"Y=C+I+G\\
Y=0.5(1-t)Y+800-30i+700\\
Y=(0.5-0.5t)Y+1500-30i\\
Y=(0.5-0.1)Y+1500-30i\\
Y=0.4Y+1500-30i\\
Y-0.4Y=1500-30i\\
0.6Y=1500-30i\\
Y=2500-50i"


LM CURVE

"Md=Ms\\
600=0.3Y-60i\\
0.3Y-60i=600\\
0.3Y=600+60i\\
Y=2000+200i"






"r*=2" and "Y=2400"


IS=LM

"2500-50i=2000+200i\\
2500-2000=200+50i\\
250i=500\\
i=2\%"

EQ

"Y=2000+200i\\
Y=2000+(200\times2)\\
Y=2400"